write the equation of the parabola y=x^2-18x-76 in standard form.
a. y-9=(x-9)^2 c. y+76=(x-9)^2
b. y+81=(x-9)^2 d. y+157=(x-9)^2

Answer by lwsshak3(11628)

write the equation of the parabola y=x^2-18x-76 in standard form.
y=x^2-18x-76
complete the square:
y=(x^2-18x+81)-81-76
y=(x-9)^2-157
(y+157)=(x-9)^2 (ans. d)
This is a parabola that opens up with vertex at (9,-157)